n - (- 3 7/10) = 1 2/3
Answer by ewatrrr(24785)

 
Hi,
n - (- 3 7/10) = 1 2/3
n + 37/10 = 5/3
n = 5/3 - 37/10 = 50/30 - 111/30 = 61/30 = 2 130